# Towards Energy-Efficient Distributed Agreement

### v1

@@ 26      $Y\leftarrow 1$
We now focus on the agreement property: The cases where every player starts with the same input value are already covered by Lemma 4.2, and henceforth we can assume that there are players starting with 00 as well as with $1$.
In our analysis, we keep track of a certain subset of the players that have adopted $1$ as their value. To this end, we define a set $P_{r}$ as follows: For round $r=1$, we have $P_{r}=\{p_{1}\}$, where $p_{1}$ is the player with the largest ID who had an input value of $1$. For $r>1$, we define $P_{r}=P_{r-1}\cup R_{r-1}$, where $R_{r}$ is the set of players that have received a message with value $1$ in round $r-1$. Note that $P_{r}$ may not include every player who had $1$ as their initial input value (unless it received a message with value $1$), and some of the players in $P_{r}$ may have already crashed by round $r$. In Phase 3 of Algorithm 3, every node that has value $1$ at the start of round $f+1$ broadcasts, causing every player who receives such a message to decide on $1$. Since at most $f$ players crash, we obtain the following as a consequence from this observation:
@@
If there is a round $r\leq f+1$ such that $|P_{r}|\geq f+1$, then the algorithm reaches agreement on $1$.
In the next lemma, we show that, as long as $|P_{r}|$ is small, that $1$-valued message are not “wasted”, in the sense that they only reach players that have not yet received such a message earlier on.
@@
If $|P_{r+1}|\leq f$ at the start of round $r+1\leq f+1$, then there does not exist a player that has received a $1$-valued message in two distinct rounds $r^{\prime}$ and $r^{\prime\prime}$ such that $r^{\prime}<r^{\prime\prime}\leq r$.
Proof 4.6.
Assume towards a contradiction that there is a set of players $Q$ violating the conclusion of the lemma. That is, for each $p_{i}\in Q$, it holds that $p_{i}$ updates $Y\leftarrow 1$ upon receiving a message in some round $r^{\prime}<r$, and $p_{i}$ also receives a message with value $1$ in a later round $r^{\prime\prime}\leq r$. Let $p_{i}\in Q$ be the player that minimizes the round number $r^{\prime\prime}$ over all players in $Q$, i.e., $r^{\prime\prime}$ is the first round in the execution in which some player receives a $1$-valued message for the second time. Since $p_{i}$ was awake in round $r^{\prime\prime}$ and received a $1$-valued message, it follows that $p_{i}\in C_{r^{\prime\prime}}$. Recall that we created $f$ committees in total, whereby the members of each committee $C_{h}$ wake up in round $h$. Moreover, JoinCommittees (i.e. Algorithm 1) ensures that the indices of the committees that player $p_{i}$ is a member of are spaced at least $\left\lceil\sqrt{n}\right\rceil$ apart, and thus $r^{\prime\prime}\geq r^{\prime}+\left\lceil\sqrt{n}\right\rceil$. According to the algorithm, $p_{i}$ sends a message to every member of the currently awake committee in each round $r^{\prime}+1,\ldots,r^{\prime}+\left\lceil\frac{f+1}{\sqrt{n}}\right\rceil$, which yields a set $M$ of $\left\lceil\sqrt{n}\right\rceil\cdot\left\lceil\frac{f+1}{\sqrt{n}}\right\rceil\geq f+1$ messages in total. Since $r^{\prime\prime}$ was the earliest round in which the lemma fails, it follows that every message in $M$ sent by $p_{i}$ in some round $s<r^{\prime\prime}$ is received by a node $p_{j}$ that has not yet received a $1$-valued message before round $s$. Note that only one of these messages could have been sent to $p_{1}\in P_{1}$, which means that $p_{j}\notin P_{s}$ and $p_{j}\in P_{s+1}$, for every player $p_{j}\neq p_{1}$, who is the destination of a message in $M$. This tells us that $P_{r^{\prime\prime}}$ must contain at least $|M|-1$ additional players, and thus
$|P_{r+1}|\geq|P_{r^{\prime\prime}}|\geq|P_{1}|+(|M|-1)\geq f+1,$
which yields the sought contradiction.
Lemma 4.7.
If a player sends a $1$-valued message to all members of at least $\left\lceil\frac{{f+1}}{\sqrt{n}}\right\rceil$ committees by the end of round $f$, then $|P_{f+1}|\geq f+1$.
Proof 4.8.
Suppose that a player $p$ does not crash before executing Lines 11-13 $\left\lceil\frac{{f+1}}{\sqrt{n}}\right\rceil$ times in some round $t$. We can assume that $|P_{t}|\leq f$, as otherwise the lemma follows immediately since $P_{t}\subseteq P_{f+1}$. Considering that each committee has a size of $\left\lceil\sqrt{n}\right\rceil$, it follows that they contain at least $\left\lceil\sqrt{f+1}/n\right\rceil\cdot\left\lceil\sqrt{n}\right\rceil\geq f+1$ players in total. By Lemma 4.5, all of these players are distinct and receive a $1$-valued message for the first time. Moreover, $p$ can be a member of at most one of these committees. Thus, we have $|P_{f+1}|\geq|P_{t+1}|\geq|P_{1}|+(f+1)-1\geq f+1$, as required.
Lemma 4.9.
It holds that either $|P_{f+1}|\geq f+1$ or the algorithm reaches agreement on 00.
Proof 4.10.
Recall that $P_{1}=\left\{p_{1}\right\}$, where $p_{1}$ is a player who had input $1$. By the code of the algorithm, $p_{1}$ attempts to send $1$ to the members of the first $\left\lceil\frac{f+1}{\sqrt{n}}\right\rceil$ committees following $C_{1}$ in Phase 2, each of which has size $\geq\sqrt{n}$. Note that Phase 2 consists of rounds $2,\ldots,f$, and it holds that
$f\geq\left\lceil\frac{f+1}{\sqrt{n}}\right\rceil-1.$ (1)
Thus, there are at least $\left(\left\lceil\frac{f+1}{\sqrt{n}}\right\rceil-1\right)\cdot\left\lceil\sqrt{n}\right\rceil$ players in the committees that are targeted by $p_{1}$ during Phase 2, which, together with the members of $C_{1}$, yields at least $\left\lceil\frac{f+1}{\sqrt{n}}\right\rceil\cdot\left\lceil\sqrt{n}\right\rceil$ players that will receive a message from $p_{1}$, unless it crashes. Similarly as in the proof of Lemma 4.5, we know that the JoinCommittees procedure ensures that $p_{1}$ is a member of at most one of these committees. It follows from Lemma 4.5 that no player is a member of more than one of these committees to which $p_{1}$ is sending a message to, and thus every $p_{j}$ that receives a message from $p_{1}$ in some round $r^{\prime}$ has not received a $1$-valued message in any earlier round, which ensures that $p_{j}\in P_{r^{\prime}+1}\setminus P_{r^{\prime}}$. Thus, $|P_{r^{\prime}+1}|\geq|P_{r^{\prime}}|+1$ and, inductively,
$|P_{r^{\prime}+1}|\geq r^{\prime}+1.$ (2)
By Lemma 4.7, we can assume that no player finishes forwarding all its $1$-valued messages, which entails that $p_{1}$ must have crashed in some round $t\leq f$. Note that (2) implies $|P_{t+1}|\geq t+1$. Furthermore, for all $r^{\prime}\in[t,f+1]$, we assume towards a contradiction that $|P_{r}|\leq f$. First, suppose there exists a round $r^{\prime}\in[t,f+1]$ in which no $1$-valued message is sent. It follows that every player who had $Y=1$ or $X=1$ must have crashed prior to round $r^{\prime}$. This means that all remaining alive players have $Y=0$ and $X=0$, and hence they will reach agreement on 00.
Therefore, in the remainder of the proof, we can assume that some $1$-valued message is sent in each round $r^{\prime}\in[t,f+1]$. Note that each committee contains at least one player $p^{\prime}$ distinct from $p_{1}$ and since we assumed that $|P_{r^{\prime}}|\leq f$, Lemma 4.5 ensures that $p^{\prime}\in P_{r^{\prime}+1}\setminus P_{r^{\prime}}$. Applying this argument inductively for all rounds in $[t,f+1]$, completes the proof.
The next lemma is a straightforward consequence of combining Lemmas 4.9 and 4.4:
Lemma 4.11.
Algorithm 3 satisfies agreement.
Lemma 4.12.
@@
Algorithm 3 sends $\mathcal{O}(n^{2})$ messages of constant size and has an energy complexity of $\mathcal{O}(\lceil{f}/{\sqrt{n}}\rceil)$.
Proof 4.13.
For the message complexity, during the first phase, every player may send at most $\mathcal{O}(\sqrt{n})$ messages. During the second phase, every player may send at most $\mathcal{O}(\sqrt{n})$ messages for $\mathcal{O}(\frac{f}{\sqrt{n}})$ consecutive rounds, leading to $\mathcal{O}(f)$ messages for each player. In the last phase, each alive player sends $n-1$ messages. Overall, this yields a message complexity of $\mathcal{O}(n^{2})$.
We next analyze the energy complexity. A given player $p$ is awake in some round $r$ due to one of the following @O reasons:
  1. 1.
    Player $p$ belongs to committee $C_{r}$. As we have $f$ committees of size $\Theta\left(\sqrt{n}\right)$ to which the players are assigned to uniformly, each player belongs to at most $\mathcal{O}(\frac{f\cdot\sqrt{n}}{n})$ committees, which contributes $\mathcal{O}(\frac{f}{\sqrt{n}})$ to the number of awake rounds by that player.
  2. 2.
    Player $p$ has input value $1$ or receives a $1$-valued message before round $f+1$. In that case, $p$ stays awake for the next $\lceil\frac{f+1}{\sqrt{n}}\rceil$ consecutive rounds to forward the message to the subsequent committees. Note that this case can arise at most twice for $p$.
@@
Overall, the energy complexity for each player is $\mathcal{O}\left(\frac{f}{\sqrt{n}}+\frac{f\cdot\sqrt{n}}{n}\right)=\mathcal{O}(\frac{f}{\sqrt{n}})$.
We remark that our algorithm can work with committees of size $k$ if players stay awake upon receiving a $1$-valued message to inform the next $\frac{f}{k}$ committees. It is easy to see that $k=\sqrt{n}$ minimizes the energy complexity, which provides some intuition for our parameter choice for Algorithm 3. This procedure reaches a consensus with a better energy consumption at the cost of the binary input value restriction, while holding the bound of optimal round complexity. However, we observe a different message complexity in comparison with the multi-value consensus. This difference can be explained by the idea that during the last round, it is not clear who holds 1 and how many of those players need to be awake for ensuring agreement on 1. To address this issue, we need every player holding 1 to broadcast leading in the worst case to $\mathcal{O}(n^{2})$ messages sent in the last rounds. In comparison, in multi-value consensus, the last committee of size $f+1$ contains sufficiently many players aware of the proposed decision value, leading to a message complexity of at most $\mathcal{O}(f.n)$ in this last round.

### v2

@@ 4.2 Notation and Terminology
We now focus on the agreement property: The cases where every player starts with the same input value are already covered by Lemma 4.7, and henceforth we can assume that there are players starting with 0 as well as with $1$.
We now prove the main correctness claim, namely, that the agreement requirement is satisfied by the algorithm. Assume towards contradiction that there exist an execution $EX$ and two non-faulty players $p^{0},p^{1}$ that decide $0,1$ respectively in $EX$. By the algorithm’s description, $p^{1}$ must have received a 1-valued message during round $f+1$ (not before, otherwise by Claim 1, $p^{0}$ would decide 1). Denote by $\hat{p}_{\ell}$ the player that sent $p^{1}$ the 1 valued message during round $f+1$ (if there are more than one, then choose the one that was triggered to do so first, and if there are still more than one, choose one arbitrarily). Note that $\hat{p}_{\ell}$ must have been triggered (to send that message) by some player $\hat{p}_{\ell-1}$ during some prior round $r_{\ell-1}$ (the same rules for picking $\hat{p}_{\ell-1}$ apply). In general, $\hat{p}_{j+1}$ is either triggered by $\hat{p}_{j}$ during round $r_{j}$ or is triggered by the algorithm before round $1$. We obtain a sequence of players $\hat{p}_{1},\dots,\hat{p}_{\ell+1}$, where $\hat{p}_{1}$ is some player that has input 1 (and is triggered by the algorithm before round $1$), $\hat{p}_{\ell+1}=p^{1}$, and $r_{\ell}=f+1$. For ease of notation we will refer to $\hat{p}_{j}$ sending 1 to $\hat{p}_{j+1}$ as $\hat{p}_{j}$ triggering $\hat{p}_{j+1}$.
{observation}
For every round $\in[h+1,f+1], r=r$$r_{j}$ for some $j\in\ell$.
Proof 4.9.
We prove the observation inductively on the round number $r$.
Base ($r=f+1$): $r=f+1=r_{\ell}$ by definition.
Step $(r<f+1)$: By the induction hypothesis, we know that for $r+1=r_{j}$ for some $j\leq\ell$. Recall that $\hat{p}_{j}$ is triggered by $\hat{p}_{j-1}$ during round $r_{j-1}<r_{j}=r+1$ ($r_{j-1}\leq r$) and that $\hat{p}_{j}$ triggers $\hat{p}_{j+1}$ during round $r_{j}=r+1$. If $r>r_{j-1}$, then $\hat{p}_{j}$ does not crash during round $r$ and sends 1 to every player in $C_{r}$. By Corollary 4.5, every non-faulty player decides 1, since $|C_{r}|\geq f+1$.
{observation}
In execution $EX$, a player that is triggered in round $r<h$ cannot be triggered during round $r^{\prime}\geq h$.
Proof 4.10.
Assume towards contradiction that $p$ is triggered during both rounds $r<h$ and $r^{\prime}\geq h$. Hence, $p$ does not crash during round $h$, and since it was triggered during round $r<h$, it has $Y=1$ during round $h$, which means it sends 1 to $C_{h}$. If $h=f+1$, then $p^{0}$ receives and decides 1; otherwise, by Corollary 4.5 ($|C_{h}|=f+1$), it follows again that $p^{0}$ decides 1, which is a contradiction.
Claim 2.
Every player in the sequence $\hat{p}_{1},\dots,\hat{p}_{\ell}$ is unique.
Proof 4.11.
Assume towards a contradiction that $\hat{p}_{j}=\hat{p}_{j^{\prime}}$ for some indices $j<j^{\prime}$. Consider the following cases. If $h\leq r_{j}$ (Phase 3), then $\hat{p}_{j}=\hat{p}_{j^{\prime}}$ sends a message during round $r_{j^{\prime}}$, which means that it did not crash during round $r_{j}$ and sent 1 to every player in $C_{r_{j}+1}$. In Phase 3, the committee size is $|C_{r_{j}}|=f+1$. Therefore, by Corollary 4.5, all non-faulty players will decide 1, in contradiction.
If $r_{j}<h$ (Phase 1 or 2), then $\hat{p}_{j}$ is triggered during round $r_{j-1}$ and $\hat{p}_{j^{\prime}}$ is triggered during round $r_{j^{\prime}-1}$. Note that $r_{j-1}<r_{j}\leq r_{j^{\prime}-1}$. It cannot be that $r_{j^{\prime}-1}\geq h$ because otherwise $\hat{p}_{j}=\hat{p}_{j^{\prime}}$ is triggered in both $r_{j-1}<h$ and $r_{j^{\prime}-1}\geq h$ violating Observation 4.3. Also, it cannot be that $r_{j^{\prime}-1}<h$, because a player can be triggered only once during rounds $[1,h]$. Hence, $r_{j^{\prime}-1}$ cannot exist.
Claim 3.
Every player that sets its timer $T>0$ during execution $EX$ must have crashed before it sets $T\leftarrow 0$.
Proof 4.12.
Let $p$ be a player that sets its timer during round $r\leq f$. Consider two cases.
If $r\geq h$ and $h<f+1$ (Phase 3), then $p$ sets its timer to 1. In the following round, $r+1$, $p$ sends 1 to $C_{r+1}$. If $r+1\leq f$ then $|C_{r+1}|=f+1$ and by Corollary 4.5, every non-faulty player decides 1. If $r+1=f+1$, then $p\in C_{f}$, $|C_{r+1}|=n$ and every non-faulty player receives 1 and thus decides 1. In both cases all non-faulty players decide 1, contradiction.
If $r<h$ (Phase 1 or 2), then $p$ sets its timer to $\lceil(f+1)/\sqrt{n^{\prime}}\rceil$. If $p$ does not crash before its timer reaches 0, then it sends 1 to the committees awake during the $\lceil\frac{f+1}{\sqrt{n^{\prime}}}\rceil$ rounds following $r$. Since the size of every such committee is at least $\sqrt{n^{\prime}}$ and by Observation 4.3, $p$ sends 1 to at least $\min\{f+1,n^{\prime}\}$ different players. If $f\leq n^{\prime}-\sqrt{n}$ or $p$ sends 1 during round $h$, then $p$ sends 1 to at least $f+1$ players, and by Corollary 4.5, every non-faulty player decides 1, contradiction.
The remaining, more difficult case is when $f>n^{\prime}-\sqrt{n^{\prime}}$ and $p$ does not send 1 during round $h$. In this case, $p$ sends 1 to all of the first $n^{\prime}<f+1$ players and $h=n^{\prime}-\sqrt{n^{\prime}}+1$. All of those players must crash before round $h+1$ (otherwise, every player in $C_{h}$ receives 1 and Corollary 4.5 applies). Recall that for every round $\in[h+1,f+1], r=r$$r_{j}$ for some $j\in[1,\ell]$ (Observation 4.3) and there is a unique player $\hat{p}_{j}$ that sends 1 during round $r$ (by Claim 2). The player $\hat{p}_{j}$ cannot have been triggered during any round $r^{\prime}<h$ by Observation 4.3, so $\hat{p}_{j}\notin\{p_{1},\ldots,p_{n^{\prime}}\}$ (as they were triggered during a round in [1,h1][1,h-1]). If $\hat{p}_{j}$ is non-faulty, then its timer reaches 0 and every non-faulty player decides 1 (as explained above). Hence, the set of players $\{p_{1},p_{2},\dots,p_{n^{\prime}}\}\cup\{\hat{p}_{j}\mid r_{j}\in[h+1,f+1]\}$ contains exactly $n^{\prime}+f-h+1=f+\sqrt{n^{\prime}}\geq f+1$ unique players that must crash, contradiction.
{observation}
In every round $r$ such that $r_{j}<r<r_{j+1}$ for some $j\in[1,\ell]$, every player in the committee $C_{r}$ receives a 1-valued message.
Proof 4.13.
By definition, during round $r_{j+1}$ the player $\hat{p}_{j+1}$ sends 1 to $\hat{p}_{j+2}$ and during round $r_{j}$, $\hat{p}_{j+1}$ is triggered by $\hat{p}_{j}$. Hence, in all rounds $r_{j}<r<r_{j+1}$, $\hat{p}_{j+1}$ does not crash (otherwise it will not be able to send during round $r_{j+1}$) and must have a non-zero timer, or in other words send a message, by Claim 3. Hence, every players in $C_{r}$ receives 1 from $\hat{p}_{j+1}$.
A failure mapping for execution $EX$ is a mapping $\varphi:[0,f]\rightarrow V$ that maps every round $j\in[0,f]$ to a player $\varphi(j)$ such that
  1. (P1)
    For $j\in[1,f]$, $\varphi(j)\in C_{j}$ and $\varphi(j)$ has crashed in $EX$.
  2. (P2)
    $\varphi(0)=\hat{p}_{1}$.
  3. (P3)
    $\varphi$ is an injection, i.e., the players $\varphi(j)$ are distinct.
Note that the existence of a failure mapping for execution $EX$ leads to contradiction, since the combination of these properties means that every round $j\in[0,f]$ is associated with a unique player that crashes in $EX$ (otherwise, $p^{0}$ decides 1), which is $f+1$ unique players, in contradiction to the number of faults being $f$.
We now describe a procedure Constr that, given an execution $EX$, attempts constructing a corresponding failure mapping $\varphi$ for $EX$. Procedure $\mbox{\sf Constr}(EX)$ operates as follows:
  1. 1.
    Let $R=\{r_{j}\mid j\in[1,\ell-1]\}$ and $P=\{\hat{p}_{j}\mid j\in[1,\ell]\}$.
  2. 2.
    $\varphi(0)=\hat{p}_{1}$ (satisfies property (P2)).
  3. 3.
    For $r_{j}\in R$, $\varphi(r_{j})=\hat{p}_{j+1}$.
  4. 4.
    For $r\in[1,\dots,f]\setminus R$, we define $\varphi(r)=p\in C_{r}$, where $p$ is chosen such that there is no $r^{\prime}<r$ with $\varphi(r^{\prime})=p$.
Note that a-priori, Procedure $\mbox{\sf Constr}(EX)$ might fail to produce a failure mapping for $EX$. This might happen if at some round $r\notin R$, there is no player in $C_{r}$ that is not already used in a previous round (since the players $\hat{p}_{j}$ for $j\in[1,\ell]$ are $m$ unique players by claim 2). Note that $r<h$.
Claim 4.
If Procedure $\mbox{\sf Constr}(EX)$ is completed successfully, then Properties (P1), (P2), (P3) are satisfied.
Proof 4.14.
Property (P1) is satisfied by the construction since for every round $r_{j}\in R$, by definition, $\hat{p}_{j+1}\in C_{r_{j}}$ receives a 1-valued message, and, for every round $r\in[1,f]\setminus R$, every player $p\in C_{r}$ receives a 1-valued message, by Observation 4.3.
Property (P2) is satisfied since $\varphi(0)=\hat{p}_{1}$ by construction.
Property (P3) is satisfied because of the following argument. Let $r<r^{\prime}$ be two rounds. If $r^{\prime}\in\{1,\dots,f\}\setminus R$, then by construction $\varphi(r)\neq\varphi(r^{\prime})$. Otherwise ($r^{\prime}\in R$), if $r\in R$ then $\varphi(r)\neq\varphi(r^{\prime})$ by Claim 2. If $r\in\{1,\dots,f\}\setminus R$ and $r^{\prime}=r_{j}$ for some $j\in[1,\ell-1]$ then $\varphi(r^{\prime})=p_{j+1}$. $p_{j+1}\notin C_{r}$ since (1) $r<h$ (2) $\hat{p}_{j+1}$ receives 1 and sets its timer during round $r_{j}$ (3) if $r_{j}\geq h$, then during round $h$, $\hat{p}_{j+1}$ sends 1 to $C_{h}$ in contradiction. Otherwise $r_{j}<h$, but a player can set its timer only once during round [1,h1][1,h-1], a contradiction. Hence, $p_{j+1}\neq\varphi(r)\in C_{r}$.
To complete the proof, we argue that this does not happen, i.e., the procedure always succeeds (hence leading to contradiction as discussed earlier).
Claim 5.
Procedure $\mbox{\sf Constr}(EX)$ is always successful.
Proof 4.15.
Assume towards contradiction that the construction is unsuccessful. Hence, there exists some round $r\notin R$, such that $r\leq\min\{f,n^{\prime}-\sqrt{n^{\prime}}\}=h-1$, where there is no player in $C_{r}$ that is not already used in a previous round. Recall that $|C_{r}|=\sqrt{n^{\prime}}$. For every $p\in C_{r}$ there exists $r^{\prime}<r$ such that $\varphi(r^{\prime})=p$. Let $p^{\prime}_{1},p^{\prime}_{2},\dots p^{\prime}_{\sqrt{n^{\prime}}}\in C_{r}$ be the players in $C_{r}$ in order of round labeling, and let $r(p^{\prime}_{j})<r$ be the round such that $\varphi(r(p^{\prime}_{j}))=p^{\prime}_{j}$ for $j\in[1,\sqrt{n^{\prime}}]$. It might be the case that $p^{\prime}_{1}=\hat{p}_{1}$ in which case $r(p^{\prime}_{1})=0$. Hence, $r(p^{\prime}_{2})\geq 1$ and by Observation 4.3, $r(p^{\prime}_{j+1})\geq r(p^{\prime}_{j})+\sqrt{n^{\prime}}$. Therefore, by induction we get that $r\geq r(p^{\prime}_{\sqrt{n^{\prime}}})+\sqrt{n^{\prime}}\geq 1+(\sqrt{n^{\prime}}-1)\sqrt{n^{\prime}}=1+n^{\prime}-\sqrt{n^{\prime}}>\min\{f,n^{\prime}-\sqrt{n^{\prime}}\}=h-1$ in contradiction.
The last claim implies that the existence of execution $EX$ leads to the existence of a failure mapping for it, in contradiction with the assumption of at most $f$ failures. This yields the following.
@@
Algorithm 3 (Binary Consensus) achieves agreement.
@@
Algorithm 3 sends $\mathcal{O}(nf)$ messages of constant size and has an energy complexity of $\mathcal{O}(\lceil{f}/{\sqrt{n}}\rceil)$.
Proof 4.18.
We start by proving the message complexity. Note that if $f\leq\sqrt{n}$, then we can simply use Algorithm 2, which, in that case, already achieves awake complexity $O(1)$, round complexity $f+1$ and message complexity $\mathcal{O}(fn)$; note that the players know $f$ and $\sqrt{n}$ and hence can locally determine which algorithm to execute. Hence, in the remainder of the analysis, we can assume that $f>\sqrt{n}$.
(Phase 1):
Every player with input value $1$ may send a message to committee $C_{1}$. Therefore $\mathcal{O}(n\cdot\sqrt{n^{\prime}})=\mathcal{O}(nf)$ message might be sent since $|C_{1}|=\sqrt{n^{\prime}}$.
(Phase 2):
Every player $p$ can be triggered only once during rounds $[1,h]$ (Observation 4.2), resulting in $p$ sending messages to the current round’s committee for $\lceil\frac{f+1}{\sqrt{n^{\prime}}}\rceil$ rounds, and the size of a committee in phase 2 is $\sqrt{n^{\prime}}$. Hence, the message complexity of this phase is $\mathcal{O}(nf+n\sqrt{n})=\mathcal{O}(nf)$.
(Phase 3):
In Phase 3, every player may send messages in round $r$ to all players in $C_{r+1}$ at most once. Since every committee in Phase 3 is of size $f+1$, this accounts for at most $\mathcal{O}(nf)$ messages.
(Phase 4):
In round $f$, every player may send a message to every player in $C_{f}$. In the final round, up to $f+1$ players may broadcast, which again yields $\mathcal{O}(nf)$ messages.
Next, we analyze the energy complexity:
(Phase 1 and 4):
The first and last phase consist of a constant number of rounds, which amounts to $\mathcal{O}(1)$ awake rounds per node.
(Phase 2):
Recall that a player is awake in round $r$ if its timer is not 0 or it belongs to $C_{r}$. During the rounds $1\leq r<h$, the timer of a player $p$ can be set to $\lceil(f+1)/\sqrt{n^{\prime}}\rceil=\mathcal{O}(f/\sqrt{n})$ only once. Therefore, every player is awake in at most
@@
$O\left(\frac{\min\{f,n^{\prime}-\sqrt{n^{\prime}}\}\cdot\sqrt{n^{\prime}}}{n^{\prime}}+\left\lceil\frac{f+1}{\sqrt{n^{\prime}}}\right\rceil\right)=\mathcal{O}(f/\sqrt{n})$
rounds.
(Phase 3):
During rounds $r\geq h$, the timer of a player $p$ can be set to 1 once, and it belongs to.
$O\left(\frac{\max\{0,f-(n^{\prime}-\sqrt{n^{\prime}})\}\cdot(f+1)}{n}\right)=\mathcal{O}(f\sqrt{n}/n)=\mathcal{O}(f/\sqrt{n})$
committees, where the first equality follows since
$f-(n^{\prime}-\sqrt{n^{\prime}})\leq n-\sqrt{n^{\prime}}(\sqrt{n^{\prime}}-1)\leq n-(\sqrt{n}-1)(\sqrt{n}-2)=3\sqrt{n}-2$
Hence, the awake complexity is $\mathcal{O}(\lceil f/\sqrt{n}\rceil)$.